OpenStudy (anonymous):

what is the solution of the equation? x+1=13

OpenStudy (aravindg):

12

OpenStudy (aravindg):

u ask silly questions lol

OpenStudy (anonymous):

\[x+1=13\] \[x+1-1=13-1\] \[x=12\]

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

(: